OpenStudy (anonymous):

write the inverse of the given function A.f(x)=2x-4 B.y=2^x

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (jhannybean):

To write the inverse: a) \(f(x) = 2x - 4\) 1. write it in terms of y: \( y= 2x - 4\) 2. switch y and x :\(x=2y -4\) 3. solve for y again: \(x+4 = 2y \implies y = \frac{x+4}{2}\)

OpenStudy (anonymous):

x+4=2y⟹y=x+42 so this is the part a? what is part b?

OpenStudy (jhannybean):

try part b on your own following the same steps :)

OpenStudy (anonymous):

It's just one step really

OpenStudy (jhannybean):

I'll help you if you get stuck

OpenStudy (anonymous):

Hint: Logarithms

OpenStudy (anonymous):

so y=2^x what would i do first?